# Online Housing Market

### v1

@@ 4.1 The Ascending Departure Permutation
The ascending departure permutation $\delta$ is $SOC$.
@@
I must show that for each agent $i$, and for any positions $i^{\prime}$ and $j^{\prime}$, such that $\Pi_{i^{\prime}}(I)=i$ and $j^{\prime}\leq i^{\prime}$, the equality $\Pi_{j^{\prime}}(I)=\Pi_{j^{\prime}}(I_{<d_{i}})$ holds when $\Pi=\delta$. Note first that agents $\Pi_{1}(I^{\prime})$, $\Pi_{2}(I^{\prime})$, $\ldots$, $\Pi_{i^{\prime}}(I^{\prime})$ are the $i^{\prime}$ agents with the smallest departure times in $I^{\prime}$. Instance $I$ is an upper set of $I^{\prime}$. Furthermore, any agent $k$ that belongs to $I$ but not to $I^{\prime}$ must have an arrival time later than $d_{i}$, since $d_{k}>a_{k}>d_{i}$. Therefore, the $i^{\prime}$ agents with the smallest departure times are the same in both $I$ and $I^{\prime}$. Finally, their relative order remains consistent in both $\Pi(I)$ and $\Pi(I^{\prime})$ because their departure times are identical in both instances. ∎
@@
Let me run Algorithm 1 with the ascending departure permutation $\delta$ on the instance described in Figure 1. At the departure of agent 2, she selects item $e_{1}$ that she prefers to $e_{2}$. Then, at the departure of agent 3, she selects item $e_{2}$ that she prefers to $e_{3}$. Finally, agent 1 leaves the market with $e_{3}$, which is the only remaining item. The resulting allocation corresponds to $M^{\prime}$ described in Figure 1.
@@
The main argument of the proof is that the matching $M$ returned by Algorithm 1 must be lexicographically optimal among the matchings of $\mathcal{M}(I)$ according to the order provided by the ascending departure permutation $\delta$ applied to $I$. Let $\mathcal{P}$ denote the subset of allocations that may Pareto-dominate $M$. According to Algorithm 1, $M(\delta(1))$ must be the most favored item of agent $\delta(1)$ in $E_{<d_{\delta(1)}}$. Therefore, all matchings in $\mathcal{P}$ must allocate $M(\delta(1))$ to agent $\delta(1)$ to Pareto-dominate $M$. Provided that $M(\delta(1))$ is allocated to agent $\delta(1)$, Algorithm 1 allocates to agent $\delta(2)$ her most favored item in $E_{<d_{\delta(2)}}\setminus\{M(\delta(1))\}$. Thus, all matchings in $\mathcal{P}$ must allocate $M(\delta(2))$ to agent $\delta(2)$ to Pareto-dominate $M$. By applying the same reasoning to the remaining agents in the order given by $\delta$, I conclude that $\mathcal{P}$ can only contain a single allocation, which corresponds to $M$. Since a matching cannot Pareto-dominate itself, no matching in $\mathcal{M}(I)$ Pareto-dominates $M$. ∎
@@
By contradiction, let $\mathcal{A}$ denote an online exchange algorithm that behaves differently than Algorithm 1 using the ascending departure permutation $\delta$ as input. This means that there exists an instance $I$ and an agent $j$ such that $\mathcal{A}_{j}(I)$ does not correspond to the assignment made by Algorithm 1 with the ascending departure permutation $\delta$. In other words, agent $j$ does not receive her most favorite item among those in $E_{<d_{\delta(j)}}$ minus the items already assigned to agents leaving earlier than her. Furthermore, I assume without loss of generality that, among the set of agents in this situation, agent $j$ is the one who leaves the earliest. Let $M$ denote the allocation obtained by Algorithm 1 with the ascending departure permutation $\delta$ applied to $I$. Let $i$ denote the agent who receives $M(j)$ instead of agent $j$ in $\mathcal{A}(I)$, i.e., such that $\mathcal{A}_{i}(I)=M(j)$. By assumption, $d_{j}<d_{i}$ must hold since, otherwise, an agent arriving earlier than agent $j$ would not receive the same item as in $M$.
@@
To show the contradiction, I construct a new instance $I^{\prime}$ by adding dummy agents to $I$. For any agent $k$ leaving after $d_{j}$, let $k^{\prime}$ be a dummy agent such that $a_{k^{\prime}}=d_{j}+k\epsilon$, with $\epsilon>0$ small enough for all these agents to arrive earlier than any other arrival or departure, and $d_{k^{\prime}}$ large enough to leave after any other agent of $I$. The preferences of the agents in $I^{\prime}$ are as follows. For any agent $k$ that was already in $I$ and leaving after $d_{j}$, her most favorite item is $e_{k^{\prime}}$, followed by the items of $I$ ranked in the same order, and finally all the other items of $I^{\prime}$ ranked in arbitrary order. Dummy agent $i^{\prime}$’s preferences are $\mathcal{A}_{j}(I)\succ_{i^{\prime}}M(j)\succ_{i^{\prime}}\ldots$, where all the other items are ranked in arbitrary order. Finally, the most favorite item of any other dummy agent $k^{\prime}$ is $\mathcal{A}_{k}(I)$, and all other items are ranked arbitrarily.
@@
Note first that since $\mathcal{A}$ is online, both $\mathcal{A}_{j}(I^{\prime})=\mathcal{A}_{j}(I^{\prime}_{<d_{j}})$ and $\mathcal{A}_{j}(I_{<d_{j}})=\mathcal{A}_{j}(I)$ hold. Furthermore, since all dummy agents arrive later than $d_{j}$, $I_{<d_{j}}=I^{\prime}_{<d_{j}}$ holds, which implies with the former equalities that $\mathcal{A}_{j}(I^{\prime})=\mathcal{A}_{j}(I)$. Now concerning the assignment of the other agents, based on their preferences and since $\mathcal{A}$ is $\mathcal{M}(I^{\prime})$-$PO$, I can say that for each dummy agent $k^{\prime}$, I have $\mathcal{A}_{k^{\prime}}(I^{\prime})=\mathcal{A}_{k}(I)$, and for each regular agent $k$ leaving after $d_{j}$, I have $\mathcal{A}_{k}(I^{\prime})=e_{k^{\prime}}$. More precisely for $i^{\prime}$ I have $\mathcal{A}_{i^{\prime}}(I^{\prime})=\mathcal{A}_{i}(I)=M(j)$. Note that except for agents $i^{\prime}$ and $j$, any agent receives her most favorite item.
@@ 4.2 The Ascending Arrival Permutation
Propositions 3 and 4 tell us that when we want an online algorithm that always returns an efficient outcome, we should focus on Algorithm 1 using the ascending departure permutation $\delta$ as input. Furthermore, Proposition 1 and 2 attest that this algorithm is $WIC$. However, it is easy to check that Algorithm 1 using the ascending departure permutation $\delta$ as input is neither $a$-$IC$, $d$-$IC$ nor $IR$. Therefore, I should consider other permutation functions as input of Algorithm 1. Another candidate is the ascending arrival permutation $\alpha$, which ranks the agents by increasing arrival time i.e., such that $a_{\alpha(1)}<a_{\alpha(2)}<\ldots<a_{\alpha(n)}$. Note that this permutation corresponds to the identity because I have assumed that $a_{1}<a_{2}<\ldots<a_{n}$. The following proposition shows that the ascending arrival permutation $\alpha$ is consistent with Algorithm 1.
@@
The ascending arrival permutation $\alpha$ is $SOC$.
@@
Let me run Algorithm 1 with the ascending arrival permutation $\alpha$ on the instance described in Figure 1. At the departure of agent 2, agent 1 selects first and picks $e_{2}$, which she prefers to $e_{1}$. Then, agent 2 selects the last remaining item at $d_{2}$, which is $e_{1}$. Finally, at the departure of agent 3, she picks the last remaining item at $d_{3}$, which is $e_{3}$. The resulting allocation corresponds to $M^{\prime\prime}$ described in Example 1.
@@
Note first that the ascending arrival permutation $\alpha$ is $SOC$ according to Proposition 5, and therefore by Proposition 1, I know that no agent has an incentive to misreport her preferences. So I only need to show that no agent has an incentive to misreport her departure time. Note also that as soon as the arrival time is not modified, changing the departure time of an agent has no impact on the ranking dictated by the ascending arrival permutation $\alpha$. I am going to show that if the reported departure time of an agent, say $i$, is delayed after a new event, which may be the departure of another agent or the arrival of a new agent, then her most favorite remaining item is either the same or even better. More formally, let $I$ be an instance and $I^{\prime}$ be the same instance except that the new departure time $d^{\prime}_{i}$ of agent $i$ is delayed to be later than the arrival time $a_{j}$ or the departure time $d_{j}$ corresponding to the event occurring right after $d_{i}$ i.e., either (i) $d_{i}<d_{j}<d^{\prime}_{i}$ or (ii) $d_{i}<a_{j}<d^{\prime}_{i}$. In both cases (i) and (ii), all iterations of the “for” loop of Algorithm 1 applied to instance $I$ and $I^{\prime}$ before reaching $d_{i}$ are the same, since only the departure time of agent $i$ has changed. Note that if agent $i$ has to choose during one of these iterations for $I$, agent $i$ chooses the same item for $I^{\prime}$ and is ultimately assigned the same item. Therefore, I can focus on the case where agent $i$ does not choose until it leaves the market at $d_{i}$ in $I$.
@@
Let me first consider case (i), where agent $j$ leaves earlier than $i$ in $I^{\prime}$ whereas she left later in $I$. I consider at the same time the two iterations of the “for” loop corresponding to $d_{i}$ and $d_{j}$ in $I$, and $d_{j}$ and $d^{\prime}_{i}$ in $I^{\prime}$, which are the first two iterations that differ between $I$ and $I^{\prime}$. Note that since there is no event occurring between $d_{i}$ and $d_{j}$ in $I$, and between $d_{j}$ and $d^{\prime}_{i}$ in $I^{\prime}$, at the end of these two iterations, the same subset of agents picks an item for $I$ and $I^{\prime}$, which should be a subset of agents ranked before either agent $i$ or $j$ in the ascending arrival permutation $\alpha$. Furthermore, they choose their items according to the same order (the one dictated by the ascending arrival permutation $\alpha$), and from the same subset of items. Therefore, each one of them chooses the exact same item for both $I$ and $I^{\prime}$, including agent $i$ who has no incentive to delay her arrival time.
@@
Finally, let me consider case (ii), where $d_{i}<a_{j}<d^{\prime}_{i}$. Note that a new item arrives to the market, namely $e_{j}$, before agent $i$ leaves the market. This new item may have consequences on the choices made by the agents who choose before $i$ according to the ascending arrival permutation $\alpha$. More specifically, one of them may choose $e_{j}$ instead of the item they selected for $I$. Once item $e_{j}$ has been chosen, say by agent $k$, it is no longer available, but the item that agent $k$ chose in $I$ at the same phase is now available and takes the place of $e_{j}$ as an additional item to choose from. So each agent choosing before $i$ either picks the same item as in $I$, or picks $e_{j}$, or another item abandoned by an agent choosing earlier. Therefore, when it is the turn of agent $i$ to choose, the item assigned to her in $I$ is available, as well as an additional item, which may be $e_{j}$ or another item abandoned by an agent choosing earlier. Consequently, she picks either the same item as in $I$ or an item that she prefers.
@@
Let $\Pi$ be an online permutation function that is different from the ascending arrival permutation $\alpha$. This means that there is an instance $I$ with two agents $i$ and $j$, whose positions according to $\Pi(I)$ are $i^{\prime}$ and $j^{\prime}$, respectively, such that both $a_{i}>a_{j}$ and $i^{\prime}<j^{\prime}$ hold. In other words, agent $i$ arrives later but is ranked before agent $j$ according to $\Pi$. By contradiction, assume that $\Pi$ makes Algorithm 1 $d$-$IC$. Let $I^{\prime}$ be a copy of instance $I$ with the exception of the departure time $d^{\prime}_{j}$ of agent $j$, which happens earlier than $d_{j}$. More precisely, I assume that $a_{j}<d^{\prime}_{j}<a_{i}$ hold. Since Algorithm 1 applied with $\Pi$ is $d$-$IC$, the item assigned to agent $j$ for $I^{\prime}$ should be no better than the one assigned to her for $I$. Furthermore, since Algorithm 1 is online, agent $j$ should receive the exact same item for instances $I^{\prime}_{<d^{\prime}_{j}}$ and $I^{\prime}$. Assume that the preferences of the agents are such that $e_{j}$ is the most favorite item of agents $i$ and $j$, and any other agent $k$ prefers her own item $e_{k}$ to any other one. It is clear that in that case, each agent receives her own item during Algorithm 1 applied to $I^{\prime}_{<d^{\prime}_{j}}$ since each agent prefers her own item (note that agent $i$ is not part of $I^{\prime}_{<d^{\prime}_{j}}$). It implies that agent $j$ receives also $e_{j}$ when Algorithm 1 is applied to $I^{\prime}$ with the ascending arrival permutation $\alpha$ as input. Since agent $j$ receives her most favorite item for $I^{\prime}$, she should receive the exact same item in $I$. But since agent $i$ is ranked before agent $j$ in the ascending arrival permutation $\alpha$, she picks first and chooses $e_{j}$, which is also her most favorite item. Therefore, agent $i$ should also receive item $e_{j}$ for $I$, leading to a contradiction since $e_{j}$ can be assigned only once. ∎
@@
With Algorithm 1 using the ascending arrival permutation $\alpha$, I obtain the property of $d$-$IC$. It is easy to check that the property of $a$-$IC$ is not achieved with the ascending arrival permutation $\alpha$, and I wonder whether there is a permutation function that achieves this property with Algorithm 1. The answer provided by the following proposition rules out the existence of such a permutation function.
@@
Let me now consider instance $I_{<d_{2}}$, where agent 3 is suppressed. For this instance, if agent 2 is ranked first in $\Pi(I_{<d_{2}})$, then during the first iteration of the “for” loop of Algorithm 1, she chooses item $e_{1}$, which is her most favorite item. During the next iteration, where agent 1 is alone in the market until she leaves at time $d_{1}$, she picks the only remaining item, $e_{2}$. However, since agent 1 prefers item $e_{1}$, she has an incentive to misreport her true arrival time and declare instead $a^{\prime}_{1}$, such that $d_{2}<a^{\prime}_{1}<d_{1}$, in order to retain item $e_{1}$. Therefore, since Algorithm 1 using $\Pi$ is $a$-$IC$, permutation function $\Pi(I_{<d_{2}})$ should be such that agent 1 chooses first. Hence, $\Pi_{1}(I_{<d_{2}})=1$ and $\Pi_{2}(I_{<d_{2}})=2$ should hold.
But since permutation function $\Pi$ is online, I should have $\Pi_{2}(I)=\Pi_{2}(I_{<d_{2}})=2$ and $\Pi_{1}(I)=\Pi_{1}(I_{<d_{2}})=1$. Now, consider the application of Algorithm 1 with $\Pi$ to $I$. During the first iteration of the “for” loop, agent 1, who is the first to choose according to $\Pi(I)$, should choose an item among $e_{1}$ and $e_{2}$, leading to a contradiction with the fact that agent 1 is assigned item $e_{3}$.

### v2

@@ 4.1 The Ascending Departure Permutation
The ascending departure permutation $\delta$ is $PS$.
@@
We must show that for each agent $i$, and for any positions $i^{\prime}$ and $j^{\prime}$, such that $\Pi_{i^{\prime}}(I)=i$ and $j^{\prime}\leq i^{\prime}$, the equality $\Pi_{j^{\prime}}(I)=\Pi_{j^{\prime}}(I_{<d_{i}})$ holds when $\Pi=\delta$. Note first that agents $\Pi_{1}(I^{\prime})$, $\Pi_{2}(I^{\prime})$, $\ldots$, $\Pi_{i^{\prime}}(I^{\prime})$ are the $i^{\prime}$ agents with the smallest departure times in $I^{\prime}$. Instance $I$ is an upper set of $I^{\prime}$. Furthermore, any agent $k$ that belongs to $I$ but not to $I^{\prime}$ must have an arrival time later than $d_{i}$, since $d_{k}>a_{k}>d_{i}$. Therefore, the $i^{\prime}$ agents with the smallest departure times are the same in both $I$ and $I^{\prime}$. Finally, their relative order remains consistent in both $\Pi(I)$ and $\Pi(I^{\prime})$ because their departure times are identical in both instances. ∎
@@
Let us run Algorithm 1 with the ascending departure permutation $\delta$ on the instance described in Figure 1. At the departure of agent 2, she selects item $e_{1}$ that she prefers to $e_{2}$. Then, at the departure of agent 3, she selects item $e_{2}$ that she prefers to $e_{3}$. Finally, agent 1 leaves the market with $e_{3}$, which is the only remaining item. The resulting allocation corresponds to $M^{\prime}$ described in Figure 1.
@@
The main argument of the proof is that the matching $M$ returned by Algorithm 1 must be lexicographically optimal among the matchings of $\mathcal{M}(I)$ according to the order provided by the ascending departure permutation $\delta$ applied to $I$. Let $\mathcal{P}$ denote the subset of allocations that may Pareto-dominate $M$. According to Algorithm 1, $M(\delta(1))$ must be the most favored item of agent $\delta(1)$ in $E_{<d_{\delta(1)}}$. Therefore, all matchings in $\mathcal{P}$ must allocate $M(\delta(1))$ to agent $\delta(1)$ to Pareto-dominate $M$. Provided that $M(\delta(1))$ is allocated to agent $\delta(1)$, Algorithm 1 allocates to agent $\delta(2)$ her most favored item in $E_{<d_{\delta(2)}}\setminus\{M(\delta(1))\}$. Thus, all matchings in $\mathcal{P}$ must allocate $M(\delta(2))$ to agent $\delta(2)$ to Pareto-dominate $M$. By applying the same reasoning to the remaining agents in the order given by $\delta$, we conclude that $\mathcal{P}$ can only contain a single allocation, which corresponds to $M$. Since a matching cannot Pareto-dominate itself, no matching in $\mathcal{M}(I)$ Pareto-dominates $M$. ∎
@@
By contradiction, let $\mathcal{A}$ denote an online exchange algorithm that behaves differently than Algorithm 1 using the ascending departure permutation $\delta$ as input. This means that there exists an instance $I$ and an agent $j$ such that $\mathcal{A}_{j}(I)$ does not correspond to the assignment made by Algorithm 1 with the ascending departure permutation $\delta$. In other words, agent $j$ does not receive her most favorite item among those in $E_{<d_{\delta(j)}}$ minus the items already assigned to agents leaving earlier than her. Furthermore, we assume without loss of generality that, among the set of agents in this situation, agent $j$ is the one who leaves the earliest. Let $M$ denote the allocation obtained by Algorithm 1 with the ascending departure permutation $\delta$ applied to $I$. Let $i$ denote the agent who receives $M(j)$ instead of agent $j$ in $\mathcal{A}(I)$, i.e., such that $\mathcal{A}_{i}(I)=M(j)$. By assumption, $d_{j}<d_{i}$ must hold since, otherwise, an agent arriving earlier than agent $j$ would not receive the same item as in $M$.
@@
To show the contradiction, we construct a new instance $I^{\prime}$ by adding dummy agents to $I$. For any agent $k$ leaving after $d_{j}$, let $k^{\prime}$ be a dummy agent such that $a_{k^{\prime}}=d_{j}+k\epsilon$, with $\epsilon>0$ small enough for all these agents to arrive earlier than any other arrival or departure, and $d_{k^{\prime}}$ large enough to leave after any other agent of $I$. The preferences of the agents in $I^{\prime}$ are as follows. For any agent $k$ that was already in $I$ and leaving after $d_{j}$, her most favorite item is $e_{k^{\prime}}$, followed by the items of $I$ ranked in the same order, and finally all the other items of $I^{\prime}$ ranked in arbitrary order. Dummy agent $i^{\prime}$’s preferences are $\mathcal{A}_{j}(I)\succ_{i^{\prime}}M(j)\succ_{i^{\prime}}\ldots$, where all the other items are ranked in arbitrary order. Finally, the most favorite item of any other dummy agent $k^{\prime}$ is $\mathcal{A}_{k}(I)$, and all other items are ranked arbitrarily.
@@
Note first that since $\mathcal{A}$ is online, both $\mathcal{A}_{j}(I^{\prime})=\mathcal{A}_{j}(I^{\prime}_{<d_{j}})$ and $\mathcal{A}_{j}(I_{<d_{j}})=\mathcal{A}_{j}(I)$ hold. Furthermore, since all dummy agents arrive later than $d_{j}$, $I_{<d_{j}}=I^{\prime}_{<d_{j}}$ holds, which implies with the former equalities that $\mathcal{A}_{j}(I^{\prime})=\mathcal{A}_{j}(I)$. Now concerning the assignment of the other agents, based on their preferences and since $\mathcal{A}$ is $\mathcal{M}(I^{\prime})$-$PO$, we can say that for each dummy agent $k^{\prime}$, we have $\mathcal{A}_{k^{\prime}}(I^{\prime})=\mathcal{A}_{k}(I)$, and for each regular agent $k$ leaving after $d_{j}$, we have $\mathcal{A}_{k}(I^{\prime})=e_{k^{\prime}}$. More precisely for $i^{\prime}$ we have $\mathcal{A}_{i^{\prime}}(I^{\prime})=\mathcal{A}_{i}(I)=M(j)$. Note that except for agents $i^{\prime}$ and $j$, any agent receives her most favorite item.
@@ 4.2 The Ascending Arrival Permutation
Propositions 3 and 4 tell us that when we want an online algorithm that always returns an efficient outcome, we should focus on Algorithm 1 using the ascending departure permutation $\delta$ as input. Furthermore, Proposition 1 and 2 attest that this algorithm is $WIC$. However, it is easy to check that Algorithm 1 using the ascending departure permutation $\delta$ as input is neither $a$-$IC$, $d$-$IC$ nor $IR$ [^6] Proposition 7 shows that the only permutation function that makes Algorithm 1 $d$-IC differs from $\delta$. Furthermore, Proposition 8 states that no permutation function makes Algorithm 1 $a$-IC. Finally, Example 5 presents an instance where the matching returned by Algorithm 1 using $\delta$ as input is not $IR$. Therefore, we should consider other permutation functions as input of Algorithm 1. Another candidate is the ascending arrival permutation $\alpha$, which ranks the agents by increasing arrival time i.e., such that $a_{\alpha(1)}<a_{\alpha(2)}<\ldots<a_{\alpha(n)}$. Note that this permutation corresponds to the identity because we have assumed that $a_{1}<a_{2}<\ldots<a_{n}$. The following proposition shows that the ascending arrival permutation $\alpha$ is prefix-stable.
@@
The ascending arrival permutation $\alpha$ is $PS$.
@@
Let us run Algorithm 1 with the ascending arrival permutation $\alpha$ on the instance described in Figure 1. At the departure of agent 2, agent 1 selects first and picks $e_{2}$, which she prefers to $e_{1}$. Then, agent 2 selects the last remaining item at $d_{2}$, which is $e_{1}$. Finally, at the departure of agent 3, she picks the last remaining item at $d_{3}$, which is $e_{3}$. The resulting allocation corresponds to $M^{\prime\prime}$ described in Example 1.
@@
Note first that the ascending arrival permutation $\alpha$ is $PS$ according to Proposition 5, and therefore by Proposition 1, we know that no agent has an incentive to misreport her preferences. So we only need to show that no agent has an incentive to misreport her departure time. Note also that as soon as the arrival time is not modified, changing the departure time of an agent has no impact on the ranking dictated by the ascending arrival permutation $\alpha$. We are going to show that if the reported departure time of an agent, say $i$, is delayed after a new event, which may be the departure of another agent or the arrival of a new agent, then her most favorite remaining item is either the same or even better. More formally, let $I$ be an instance and $I^{\prime}$ be the same instance except that the new departure time $d^{\prime}_{i}$ of agent $i$ is delayed to be later than the arrival time $a_{j}$ or the departure time $d_{j}$ corresponding to the event occurring right after $d_{i}$ i.e., either (i) $d_{i}<d_{j}<d^{\prime}_{i}$ or (ii) $d_{i}<a_{j}<d^{\prime}_{i}$. In both cases (i) and (ii), all iterations of the “for” loop of Algorithm 1 applied to instance $I$ and $I^{\prime}$ before reaching $d_{i}$ are the same, since only the departure time of agent $i$ has changed. Note that if agent $i$ has to choose during one of these iterations for $I$, agent $i$ chooses the same item for $I^{\prime}$ and is ultimately assigned the same item. Therefore, we can focus on the case where agent $i$ does not choose until it leaves the market at $d_{i}$ in $I$.
@@
Let us first consider case (i), where agent $j$ leaves earlier than $i$ in $I^{\prime}$ whereas she left later in $I$. We consider at the same time the two iterations of the “for” loop corresponding to $d_{i}$ and $d_{j}$ in $I$, and $d_{j}$ and $d^{\prime}_{i}$ in $I^{\prime}$, which are the first two iterations that differ between $I$ and $I^{\prime}$. Note that since there is no event occurring between $d_{i}$ and $d_{j}$ in $I$, and between $d_{j}$ and $d^{\prime}_{i}$ in $I^{\prime}$, at the end of these two iterations, the same subset of agents picks an item for $I$ and $I^{\prime}$, which should be a subset of agents ranked before either agent $i$ or $j$ in the ascending arrival permutation $\alpha$. Furthermore, they choose their items according to the same order (the one dictated by the ascending arrival permutation $\alpha$), and from the same subset of items. Therefore, each one of them chooses the exact same item for both $I$ and $I^{\prime}$, including agent $i$ who has no incentive to delay her arrival time.
@@
Finally, let us consider case (ii), where $d_{i}<a_{j}<d^{\prime}_{i}$. Note that a new item arrives to the market, namely $e_{j}$, before agent $i$ leaves the market. This new item may have consequences on the choices made by the agents who choose before $i$ according to the ascending arrival permutation $\alpha$. More specifically, one of them may choose $e_{j}$ instead of the item they selected for $I$. Once item $e_{j}$ has been chosen, say by agent $k$, it is no longer available, but the item that agent $k$ chose in $I$ at the same phase is now available and takes the place of $e_{j}$ as an additional item to choose from. So each agent choosing before $i$ either picks the same item as in $I$, or picks $e_{j}$, or another item abandoned by an agent choosing earlier. Therefore, when it is the turn of agent $i$ to choose, the item assigned to her in $I$ is available, as well as an additional item, which may be $e_{j}$ or another item abandoned by an agent choosing earlier. Consequently, she picks either the same item as in $I$ or an item that she prefers.
@@
Let $\Pi$ be an online permutation function that is different from the ascending arrival permutation $\alpha$. This means that there is an instance $I$ with two agents $i$ and $j$, whose positions according to $\Pi(I)$ are $i^{\prime}$ and $j^{\prime}$, respectively, such that both $a_{i}>a_{j}$ and $i^{\prime}<j^{\prime}$ hold. In other words, agent $i$ arrives later but is ranked before agent $j$ according to $\Pi$. By contradiction, assume that $\Pi$ makes Algorithm 1 $d$-$IC$. Let $I^{\prime}$ be a copy of instance $I$ with the exception of the departure time $d^{\prime}_{j}$ of agent $j$, which happens earlier than $d_{j}$. More precisely, we assume that $a_{j}<d^{\prime}_{j}<a_{i}$ hold. Since Algorithm 1 applied with $\Pi$ is $d$-$IC$, the item assigned to agent $j$ for $I^{\prime}$ should be no better than the one assigned to her for $I$. Furthermore, since Algorithm 1 is online, agent $j$ should receive the exact same item for instances $I^{\prime}_{<d^{\prime}_{j}}$ and $I^{\prime}$. Assume that the preferences of the agents are such that $e_{j}$ is the most favorite item of agents $i$ and $j$, and any other agent $k$ prefers her own item $e_{k}$ to any other one. It is clear that in that case, each agent receives her own item during Algorithm 1 applied to $I^{\prime}_{<d^{\prime}_{j}}$ since each agent prefers her own item (note that agent $i$ is not part of $I^{\prime}_{<d^{\prime}_{j}}$). It implies that agent $j$ receives also $e_{j}$ when Algorithm 1 is applied to $I^{\prime}$ with the ascending arrival permutation $\alpha$ as input. Since agent $j$ receives her most favorite item for $I^{\prime}$, she should receive the exact same item in $I$. But since agent $i$ is ranked before agent $j$ in the ascending arrival permutation $\alpha$, she picks first and chooses $e_{j}$, which is also her most favorite item. Therefore, agent $i$ should also receive item $e_{j}$ for $I$, leading to a contradiction since $e_{j}$ can be assigned only once. ∎
@@
With Algorithm 1 using the ascending arrival permutation $\alpha$, we obtain the property of $d$-$IC$. It is easy to check that the property of $a$-$IC$ is not achieved with the ascending arrival permutation $\alpha$, [^7] This is in fact a direct consequence of Proposition 8. and we wonder whether there is a permutation function that achieves this property with Algorithm 1. The answer provided by the following proposition rules out the existence of such a permutation function.
@@
Let us now consider instance $I_{<d_{2}}$, where agent 3 is suppressed. For this instance, if agent 2 is ranked first in $\Pi(I_{<d_{2}})$, then during the first iteration of the “for” loop of Algorithm 1, she chooses item $e_{1}$, which is her most favorite item. During the next iteration, where agent 1 is alone in the market until she leaves at time $d_{1}$, she picks the only remaining item, $e_{2}$. However, since agent 1 prefers item $e_{1}$, she has an incentive to misreport her true arrival time and declare instead $a^{\prime}_{1}$, such that $d_{2}<a^{\prime}_{1}<d_{1}$, in order to retain item $e_{1}$. Therefore, since Algorithm 1 using $\Pi$ is $a$-$IC$, permutation function $\Pi(I_{<d_{2}})$ should be such that agent 1 chooses first. Hence, $\Pi_{1}(I_{<d_{2}})=1$ and $\Pi_{2}(I_{<d_{2}})=2$ should hold.
But since permutation function $\Pi$ is online, we should have $\Pi_{2}(I)=\Pi_{2}(I_{<d_{2}})=2$ and $\Pi_{1}(I)=\Pi_{1}(I_{<d_{2}})=1$. Now, consider the application of Algorithm 1 with $\Pi$ to $I$. During the first iteration of the “for” loop, agent 1, who is the first to choose according to $\Pi(I)$, should choose an item among $e_{1}$ and $e_{2}$, leading to a contradiction with the fact that agent 1 is assigned item $e_{3}$.